OpenStudy (anonymous):

hey! you should help. because you should. Eliminate the parameter. x = 7t, y = t + 4

jimthompson5910 (jim_thompson5910):

solve for t in y = t + 4 to get y = t + 4 y-4 = t t = y-4

jimthompson5910 (jim_thompson5910):

Then plug this into the other equation x = 7t x = 7(y-4) x = 7y - 28 x - 7y = -28

OpenStudy (anonymous):

Thank you so much. that really helped! :3

jimthompson5910 (jim_thompson5910):

you're welcome